OpenStudy (anonymous):

Miley’s dog, Ella loves to play outside. Ella’s 15 ft leash is tied to a stake in the ground. How much area does she have to play?

OpenStudy (anonymous):

pi * 15^2 = 225 * pi =~ 706.858 ft^2 22 / pi =~ 7" 11.5 * 2 * pi = 23 * pi =~ 72.2566' =~ 72 feet 3 inches 4 / 2 = 2" 15 * 2 = 30 cm

OpenStudy (anonymous):

these aew the only options though: 807.5 ft2 78.5 ft2 47.1 ft2 706.9 ft2

OpenStudy (anonymous):

706.9